# Quantum computing inspired paintings: reinterpreting classical masterpieces

### v1

@@ I Introduction
Our choice of paintings follows a line from the Naturalism of Caravaggio to the Surrealism of Magritte and, lastly, the Abstraction of Richter. We consider our work from various angles. The first is to explore, how quantum computing renders the original paintings from the three epochs differently leading to new versions of the paintings. The second is the aesthetical reception of the digital images produced by the authors. And, the third is the human intervention by reproducing the digital template on a wooden panel with oil, giving it an additional human element. Finally, through using a quantum computer, and thus employing the quantum mechanical principles, we are able to represent a contemporary concept, i.e. how reality is losing its form and becoming abstract and evanescent. We titled the paintings series as “Quantum Transformation I, II and III [^1] Note that we have added a watermark in the pictures of these paintings presented here..
The paper is structured as follows: In Sec. II, we introduce the Hamiltonian of the Ising model and the time evolution via a Trotterization. We define the quantum circuit used for the computation with its parameterized gates (i.e. the building blocks of quantum circuits). The results are reported in Sec. III, for the painting “Quantum Transformation I: Caravaggio”, Sec. IV for the painting “Quantum Transformation II: Magritte”, and Sec. V for the painting “Quantum Transformation III: Richter”. In each section, we describe each of the three paintings, with the results from the quantum computation, and show the final pictures of the oil paintings [^2] The painting “Quantum Transformation II: Magritte” is still in production and will be included in a forthcoming version of this paper. We reported here the digital image with the quantum computing results.. The source code used to work on these paintings is available in Ref. [16]. To conclude the paper, Sec. VI gives a summary and the outlook for this project.
@@ III Quantum Transformation I: Caravaggio
    Therefore, there reordered list would be $\mathbf{\bar{i}}=[0,1,3,12]=[i_{0},i_{1},i_{3},i_{2}]$. As a result, the two rightmost pixels are swapped, as summarized in Fig. 4.
@@
For the quantum computation, we use the ibm_kyoto quantum hardware with $4096$ shots (measurements of the quantum circuit) and we translate the quantum computing results into new positions of the squares in the grid. Fig. 6 shows the photograph of the final oil painting, hand-painted on a wooden panel [^3] This painting will be exhibited at the European Quantum Technologies Conference (EQTC) in Lisbon (Portugal) [22]..
@@ IV Quantum Transformation II: Magritte
Figure 7: The painting “Quantum Transformation II: Magritte” (digital image): The entire picture is modified by the quantum time evolution. The only element that remains untouched is the green apple.
Also in this painting, we refer to the concept of observation and subsequent interaction with a quantum system. In our quantum aesthetics version, we apply the quantum process to the entire painting (again, as the conceptual surface of reflection) displacing all of its elements, with the exception of the apple in front of the man’s face. With this process of displacement, we are led to imagine the features of the face that now become unveiled. To create the man’s face, we edit the digital image of the painting by taking inspiration from another painting by Magritte, “La bonne foi” [24]. To reach this result, we divide the painting into 16 columns and 20 rows. We use a similar approach as in Sec. III, with the computation of Eq. (6) and the subsequent reordering of the elements. Since the computation can be quite demanding, we decided to mirror the time evolution with 10 steps, so we can reduce the number of Trotter steps and the depth of the circuits. In result, the painting’s upper half utilizes the regular time evolution (going from the first time-step in the bottom and reaching the middle at step 10), whereas the bottom half is mirrored (the middle section starting from step 10, down to one in the last row). Consequently, the middle section of the painting contained the most time-evolved content. Furthermore, instead of duplicating a single 10-step result for both halves, the final computation counted with two independent runs of the same circuit, both simulating 10 Trotter steps. The upper half was computed with ibm_sherbrooke, and the bottom one with ibm_strasbourg; both with 4096 shots per circuit. Fig. 7 illustrates the final outcome [^4] The painting is still in production and will be included in a forthcoming version of this paper., after the application of the quantum computation.
@@ V Quantum Transformation III: Richter
Figure 8: The painting “Quantum Transformation III: Richter”: (panel (a)) The original version titled “192 Farben” (digital reconstruction): The image has been produced digitally based on the original painting. (panel (b)) The revisited version (oil on wooden panel): This photograph of the oil painting illustrates the colors changed by translating the results from quantum computation. Panel size: $75\times 100$ cm.
@@
Differently from Sec. III and Sec. IV, the index $n$ is not used in the reordering, as the intention is to produce a more intense rearrangement of the colors across the entire painting (see the explanation in Sec. III). The initial state |$\ket{\psi(0)}$ is prepared by single $R_{y}(\theta_{n})$ rotation gates to let $C_{n}(0)=n$. After computing the color indices for all the time slices and sites, we reorder these 192 values from smallest to largest. The corresponding colors are then rearranged from left to right and top to bottom, following the new order of $C_{n}(t)$. As the initial color indices at $t=0$ are relatively small, they remain in the original position, i.e. the bottom row of the figure, and other color squares are reordered based on $C_{n}(t)$. We use the ibm_nazca quantum hardware for the quantum computation and again 4096 shots. After the time evolution, with Eq. (4), the final resulting artwork depicts the same colors of the original composition (Fig. 8(a)) but reordered with quantum results (Fig. 8(b)) [^5] Note that possible imperfections in the real physical painting are due to the challenge of exactly reproducing the colors by hand, as well as the difficulty of exactly reproducing the original colors in RGB..

### v4

@@ I Introduction
Our choice of paintings follows a line from the Naturalism of Caravaggio to the Surrealism of Magritte and, lastly, the Abstraction of Richter. We consider our work from various angles. The first is to explore, how quantum computing renders the original paintings from the three epochs differently leading to new versions of the paintings. The second is the aesthetical reception of the digital images produced by the authors. And, the third is the human intervention by reproducing the digital template on a wooden panel with oil, giving it an additional human element. Finally, through using a quantum computer, and thus employing the quantum mechanical principles, we are able to represent a contemporary concept, i.e. how reality is losing its form and becoming abstract and evanescent. We titled the paintings series as “Quantum Transformation I, II and III [^1] The paintings will be exhibited at the Ars Electronica Festival 2025 [16]..
The paper is structured as follows: In Sec. II, we introduce the Hamiltonian of the Ising model and the time evolution via a Trotterization. We define the quantum circuit used for the computation with its parameterized gates (i.e. the building blocks of quantum circuits). The results are reported in Sec. III, for the painting “Quantum Transformation I: Caravaggio”, Sec. IV for the painting “Quantum Transformation II: Magritte”, and Sec. V for the painting “Quantum Transformation III: Richter”. In each section, we describe each of the three paintings, with the results from the quantum computation, and show the final pictures of the oil paintings. The source code used to work on these paintings is available in Ref. [17]. To conclude the paper, Sec. VI gives a summary and the outlook for this project.
@@ III Quantum Transformation I: Caravaggio
    Therefore, the reordered list would be $\mathbf{\bar{i}}=[0,1,3,12]=[i_{0},i_{1},i_{3},i_{2}]$. As a result, the two rightmost pixels are swapped, as summarized in Fig. 4.
@@
For the quantum computation, we use the ibm_kyoto quantum hardware with $4096$ shots (measurements of the quantum circuit) and we translate the quantum computing results into new positions of the squares in the grid. Fig. 6 shows the photograph of the final oil painting, hand-painted on a wooden panel [^2] This painting has been exhibited at the European Quantum Technologies Conference (EQTC) in Lisbon (Portugal) [23]..
@@ IV Quantum Transformation II: Magritte
Also in this painting, we refer to the concept of observation and subsequent interaction with a quantum system. In our quantum aesthetics version, we apply the quantum process to the entire painting (again, as the conceptual surface of reflection) displacing all of its elements, with the exception of the apple in front of the man’s face. With this process of displacement, we are led to imagine the features of the face that now become unveiled. To create the man’s face, we edit the digital image of the painting by taking inspiration from another painting by Magritte, “La bonne foi” [25]. To reach this result, we divide the painting into 16 columns and 20 rows. We use a similar approach as in Sec. III, with the computation of Eq. (6) and the subsequent reordering of the elements. Since the computation can be quite demanding, we decided to mirror the time evolution with 10 steps, so we can reduce the number of Trotter steps and the depth of the circuits. In result, the painting’s upper half utilizes the regular time evolution (going from the first time-step in the bottom and reaching the middle at step 10), whereas the bottom half is mirrored (the middle section starting from step 10, down to one in the last row). Consequently, the middle section of the painting contained the most time-evolved content. Furthermore, instead of duplicating a single 10-step result for both halves, the final computation counted with two independent runs of the same circuit, both simulating 10 Trotter steps. The upper half was computed with ibm_sherbrooke, and the bottom one with ibm_strasbourg; both with 4096 shots per circuit. Fig. 7 illustrates the final outcome, after the application of the quantum computation.
Figure 7: The painting “Quantum Transformation II: Magritte” (oil on wooden panel): In this photograph of the oil painting the entire subject is modified by the quantum time evolution. The only element that remains untouched is the green apple. Panel size: $60\times 77$ cm.
@@ V Quantum Transformation III: Richter
Figure 8: The painting “Quantum Transformation III: Richter” (oil on wooden panel): This photograph of the oil painting illustrates the colors changed by translating the results from quantum computation. Panel size: $75\times 100$ cm.
@@
Differently from Sec. III and Sec. IV, the index $n$ is not used in the reordering, as the intention is to produce a more intense rearrangement of the colors across the entire painting (see the explanation in Sec. III). The initial state |$\ket{\psi(0)}$ is prepared by single $R_{y}(\theta_{n})$ rotation gates to let $C_{n}(0)=n$. After computing the color indices for all the time slices and sites, we reorder these 192 values from smallest to largest. The corresponding colors are then rearranged from left to right and top to bottom, following the new order of $C_{n}(t)$. As the initial color indices at $t=0$ are relatively small, they remain in the original position, i.e. the bottom row of the figure, and other color squares are reordered based on $C_{n}(t)$. We use the ibm_nazca quantum hardware for the quantum computation and again 4096 shots. After the time evolution, with Eq. (4), the final resulting artwork depicts the same colors of the original composition but reordered with quantum results (Fig. 8) [^3] Note that possible imperfections in the real physical painting are due to the challenge of exactly reproducing the colors by hand, as well as the difficulty of exactly reproducing the original colors in RGB..